OpenStudy (anonymous):

how does the 'k' from rate equation be dependent on temperature and catalyst?

OpenStudy (anonymous):

k is only depenedent on temp not on catalyst as when we added catalyst it increases in the same amount in both reactant and product concentration so it ihas no affect on k but by changing the temp the reaction can also be change by increasing temp it will favour backward or forward in the reaction through which k also be change

OpenStudy (anonymous):

k IS affected by the catalyst. If not how would the rate of reaction change? Anyway according to the Arrhenius equation, k is dependent on the activation energy. Catalysts provide a pathway with less activation energy, which will increase k.

OpenStudy (anonymous):

no its not sir it only increases the rate of reation means if u add catalyst it will increasse in both reactant and product concteration for example reatant concentration is 50 and product concentration be 10 k=10/50= 1/5 now by adding catalyst as amount5 now reactant concentration will become 250 and product will become 50 k=50/250 =1/5 only the rate of reation increaes or decreases but k will never change because it only depends on temperature!!!!!!!!!!!!!!!

OpenStudy (anonymous):

uhm sorry I'm very sure I'm correct please check your info. Arrhenius equation is dependent on Activation Energy. The catalyst provides an alternative pathway for the reaction to occur with a lower E_A. so. k is thus increased (since the exponential has a negative sign)